# Single-file dynamics with general charge measures

### 25 Jan 2024 v1

@@ 5 Conclusion and discussion
Another question concerns the development of a hydrodynamic framework for typical fluctuations in integrable systems. While ballistic macroscopic fluctuation theory [6] and space-time duality [7] give access to large fluctuations, the central limit property critically hinges on regularity of the full-counting statistics [8], which has been found to be violated in certain integrable models at half-filling that exhibit sub-ballistic transport [10, 11]. It appears plausible that sub-ballistic transport in integrable models is fundamentally related to broken regularity of the full-counting statistics, opening the way for diverse non-Gaussian fluctuations.
@@ Appendix B Typical fluctuations
    Typical fluctuations are primarily governed by the large-scale behavior of the measure which corresponds to small $k$-behavior of the Fourier transformed measure (19)
@@ B.1 Nonequilibrium typical conditional probability
    where $\overline{\nu}=-{\rm sgn}(j_{p})$. Making a change of variable $u=kt^{\alpha/2}$, the integral becomes
@@
    $\lim_{t\to\infty}\left[1+a_{1}t^{-\alpha/2}+a_{2}t^{-\alpha}\right]^{b_{1}t^{\alpha}+b_{2}t^{\alpha/2}}e^{-c_{1}t^{\alpha/2}}=\begin{cases}e^{-\frac{a_{1}^{2}b2}{2}+a_{2}b1+a_{1}b_{2}}&{\rm for}\enspace c_{1}=a_{1}b_{1},\\ \infty&{\rm for}\enspace c_{1}\neq a_{1}b_{1},\end{cases}$ (93)
@@ B.2 Equilibrium typical conditional probability
    With a change of variables $u=kt^{\alpha/4}$ and insertion of the Taylor expansion (90) the integral becomes
@@ C.1.2 Saddle point localization
    Since $\hat{\omega}_{\mu\pm}$ are entire, we can deform the integration contour in (100) without changing the value of the integrals (100). We accordingly deform the path of integration to the steepest descent contours $\mathcal{I_{\pm}}$ across the saddle points, resulting in an integral whose exponential contribution is easily evaluated as $t\to\infty$
@@ C.2 Generic properties of the split rate functions
      where positivity follows from (107) using the identification (103) and noting that a derivative with respect to $k$ gives an additional ${\rm i}$ factor. The Hessian determinant then reads

### 01 Feb 2024 v2

@@ 5 Conclusion and discussion
Another question concerns the development of a hydrodynamic framework for typical fluctuations in integrable systems. While ballistic macroscopic fluctuation theory [7] and space-time duality [8] give access to large fluctuations, the central limit property critically hinges on regularity of the full-counting statistics [9], which has been found to be violated in certain integrable models that exhibit sub-ballistic transport at half-filling [11, 12]. It appears plausible that sub-ballistic transport in integrable models is fundamentally related to broken regularity of the full-counting statistics, opening the way for diverse non-Gaussian fluctuations.
@@ Appendix B Typical fluctuations
    Typical fluctuations are primarily governed by the large-scale structure of the measure which corresponds to small $k$-behavior of the Fourier transformed measure (19)
@@ B.1 Nonequilibrium typical conditional probability
    where $\overline{\nu}=-{\rm sgn}(j_{p})$. With a change of variable $u=kt^{\alpha/2}$, the integral becomes
@@
    $\lim_{t\to\infty}\left[1+a_{1}t^{-\alpha/2}+a_{2}t^{-\alpha}\right]^{b_{1}t^{\alpha}+b_{2}t^{\alpha/2}}e^{-c_{1}t^{\alpha/2}}=\begin{cases}e^{-\frac{a_{1}^{2}b2}{2}+a_{2}b_{1}+a_{1}b_{2}}&{\rm for}\enspace c_{1}=a_{1}b_{1},\\ \infty&{\rm for}\enspace c_{1}\neq a_{1}b_{1},\end{cases}$ (95)
@@ B.2 Equilibrium typical conditional probability
    With a change of variables $u=kt^{\alpha/4}$ and insertion of the Taylor expansion (92), the integral becomes
@@ C.1.2 Saddle point localization
    Since $\hat{\omega}_{\mu_{\pm}}$ are entire, we can deform the integration contour in (102) without changing the value of the integrals (102). We accordingly deform the path of integration to the steepest descent contours $\mathcal{I_{\pm}}$ across the saddle points, resulting in an integral whose exponential contribution is easily evaluated as $t\to\infty$
@@ C.2 Generic properties of the split rate functions
      where positivity follows from (109) using the identification (105) and noting that a derivative with respect to $k$ gives an additional factor of ${\rm i}$. The Hessian determinant then reads